# FrameOracle: Learning What to See and How Much to See in Videos

### v1

@@ 4.2 Training
We train FrameOracle using a curriculum-based, four-stage protocol. This strategy progressively refines the policy $\Pi_{\theta}$, teaching it to reason effectively over the pre-sampled frames (e.g., 16 or 64). The staged training leverages four widely used public VideoQA datasets, covering clips ranging from roughly 10 seconds to 15 minutes in length. Details of the full dataset composition are provided in Appendix A.
Stage 1: Text-Visual Alignment. The initial stage focuses on learning a robust cross-modal representation by aligning the textual prompt with the visual content of the candidate frames. We use the pre-trained text-visual model SigLIP (Zhai et al., 2023) as a teacher to provide supervision. For each prompt-frame pair, a SigLIP similarity score serves as the target relevance signal. The feature projectors and the cross-modal Transformer encoder are trained with a RankNet loss (Burges et al., 2005), encouraging the model’s predicted scores to match the relative ordering of the SigLIP similarities. Concretely, for any two frames $i$ and $j$, let $s_{i}$ and $s_{j}$ denote their SigLIP similarity scores, and $y_{i}$ and $y_{j}$ their predicted scores. We define the pairwise preference label as $t_{ij}=\mathrm{sign}(s_{i}-s_{j})$. The RankNet loss is then given by
@@
where $t_{ij}=0$ corresponds to tied frames and does not contribute to the gradient. In this way, the alignment capability of SigLIP is distilled into FrameOracle. The K Head remains frozen during this stage.
Stage 2: Rank Head Optimization. In the second stage, we train the Rank Head to identify the most salient frames in the candidate set $V_{C}$. Unlike the first stage, where SigLIP-based supervision is computed independently for each frame and provides no temporal guidance, this stage uses the downstream VLM’s loss as a supervisory signal, allowing the selector to capture temporal dependencies across frames. To generate training targets, we adopt a leave-one-out (LOO) approach: for each frame $f_{i}\in V_{C}$, we remove it from the input set and pass the remaining frames through the VLM, measuring the change in its loss. A larger increase indicates that $f_{i}$ is more important. These importance scores serve as soft targets, and the Rank Head is trained with a RankNet loss to predict them. During this stage, the K Head remains frozen, while the Transformer encoder and feature projectors are fine-tuned with a smaller learning rate to stabilize training.
Stage 3: K Head Optimization. The third stage focuses on training the K Head to predict the number of frames. During this stage, the Rank Head is frozen, while the feature projectors and the Transformer encoder are fine-tuned with a very small learning rate for slight adaptation. For each sample, we evaluate the downstream VLM (the same backbone and task loss as in Stage 2) using the top-$k$ frames from $V_{C}$, ranked by the frozen Rank Head, for a candidate value $k\in N$. We then select the target
$k^{*}\;=\;\arg\min_{k\in N}\Big(\mathrm{zscore}(\mathcal{L}_{\text{task}}(k))+\lambda_{k}\,k\Big),$ (2)
where the linear penalty balances accuracy and frame cost. The K Head predicts a categorical distribution $p_{\theta}(k)$ over $k\in\{1,\dots,N\}$ and is trained with
@@
regresses the predicted expectation to $k^{*}$, and $\mathcal{L}_{\text{class}}$ is a KL divergence aligning $p_{\theta}$ with a Gaussian-shaped soft target centered at $k^{*}$.
Stage 4: Supervised Fine-tuning with Ground Truth. In the final stage, we perform supervised fine-tuning (SFT) on FrameOracle-41K, which provides supervision for both the keyframe indices and the number of frames. Unlike the earlier stages, which rely on weak or proxy signals, FrameOracle-41K offers high-quality annotations that have been verified for consistency. The Rank Head is trained to align its predictions with the annotated keyframes, while the K Head is jointly trained to match the annotated $K$ values. This strong, direct supervision further refines the selection policy beyond what is achieved in Stages 1–3.
@@ 5.1 Experiment Settings
Benchmarks. We evaluate FrameOracle on six widely adopted video benchmarks, which can be divided into long-video and short-video understanding tasks. For long-video understanding, we include EgoSchema (Mangalam et al., 2023), LongVideoBench (Wu et al., 2024), MLVU (Zhou et al., 2025), and Video-MME (Fu et al., 2025), all of which require reasoning over extended temporal contexts ranging from minutes to hours. These datasets emphasize challenges such as cross-event reasoning, global consistency, and temporal grounding across lengthy sequences. For short-video understanding, we evaluate on NExTQA (Xiao et al., 2021), and Perception (Pătrăucean et al., 2023), which involve clips typically within tens of seconds. These benchmarks focus on fine-grained event recognition, local temporal relations, and reasoning within concise videos. We follow the LMMs-Eval library (Zhang et al., 2025b) for evaluation, and report accuracy across all benchmarks.
Table 1: FrameOracle vs. SOTA VLMs. “Frames” shows $M\rightarrow\bar{K}$: FrameOracle starts from $M$ uniformly sampled frames and reduces to an average of $\bar{K}$. Highlighted rows show the upper-bound performance of FrameOracle with larger frame inputs. LVB = LongVideoBench validation set.
@@
ShareGPT4Video-8B (Chen et al., 2024b) 16 - - - - 41.8 39.9 - 46.4 -
LLaMA-VID-7B (Li et al., 2024b) 16 - - - 44.6 - 25.9 38.5 33.2 -
VideoChat2-7B (Li et al., 2024a) 16 - - - - 39.3 39.5 63.6 44.5 -
@@
5.2 Results and Analysis
Comparisons with State-of-the-Art Models. Table 5.1 presents a comprehensive comparison of FrameOracle across two categories: (1) five state-of-the-art (SOTA) VLMs, and (2) its integration with five diverse VLMs, Qwen2.5-VL (Bai et al., 2025), LLaVA-OneVision (Li et al., 2025), LLaVA-Video (Zhang et al., 2024), VideoLLaMA3 (Zhang et al., 2025a), and the proprietary GPT-4o (Hurst et al., 2024). Qwen2.5-VL internally merges every two adjacent frames into a single representation. To ensure a fair comparison with models that process raw frames directly, we report the baseline using 32 frames. For each model integrated with FrameOracle, we report results for two configurations: using 16-frame FrameOracle and 64-frame FrameOracle.
Under the 16-frame condition, FrameOracle maintains accuracy comparable to the baseline models across all benchmarks while reducing the number of frames by approximately 35%. With 64-frame inputs, FrameOracle begins with a denser candidate set and adaptively selects relevant frames. In this setting, it consistently improves performance over the baseline models while still reducing frames by about 15%. This demonstrates that a larger candidate pool enables FrameOracle to better exploit temporal redundancy, resulting in improved accuracy–efficiency trade-offs. FrameOracle is trained independently and applied in a fully plug-and-play manner, requiring no co-training or backbone-specific adaptation. These results confirm its ability to generalize across model architectures.
One might expect that providing more frames always improves performance, since additional frames offer more visual evidence. However, Table 5.1 shows the opposite: using more frames often fails to help and can even reduce accuracy. This aligns with recent findings that long-video reasoning is inherently sparse, with only a small subset of frames being truly relevant (Park et al., 2024). Extra frames primarily introduce redundancy and noise, diluting cross-modal attention and yielding diminishing returns (Li et al., 2023).
By contrast, when FrameOracle selects a smaller but more informative subset of frames, performance can improve, especially on open-ended benchmarks. For example, on LLaVA-OneVision-7B, reducing 16 frames to roughly 10.4 improves NExTQA metrics (OE_val: $14.6\rightarrow 16.1$, OE_test: $16.7\rightarrow 17.8$) and EgoSchema ($60.8\rightarrow 62.4$). Similar trends are observed for GPT-4o, where accuracy rises from 55.6 to 56.2 despite using fewer frames. Qualitative examples in Appendix E (Figure 6) further illustrate this effect: FrameOracle identifies the key evidence with fewer frames, yielding correct answers where naive higher-frame sampling fails.
Table 2: FrameOracle vs. SOTA keyframe selection methods. NExTQA reports MCQ. Methods using more frames or larger LLMs are shown in gray. LVB = LongVideoBench validation set.
@@ 5.2 Results and Analysis
We compare FrameOracle with (1) keyframe selection methods that are jointly trained with their VLM backbone in their original configurations, and (2) plug-and-play keyframe selection methods applied to open-source models (Table 5.2). The first category cannot be applied directly to open-source models, and FFS (Buch et al., 2025) is the only method that adaptively determines the number of retained frames; all other methods assume a fixed number of keyframes. All plug-and-play selection methods only provided 8-frame results. To ensure a fair comparison, we disable FrameOracle’s K Head and rely solely on the Rank Head: given 64 uniformly sampled frames, we select the top-8 ranked frames as input to the backbone VLMs.
FrameOracle achieves competitive performance compared to prior plug-and-play methods. Across NExTQA, LongVideoBench, Video-MME, and EgoSchema, it improves accuracy by roughly 2–4 percentage points, outperforming Frame-Voyager (Yu et al., 2025), BOLT (Liu et al., 2025), and KFC (Fang et al., 2025). On MLVU, FrameOracle outperforms the base VLMs but does not surpass heuristic methods such as KFC, a greedy selection strategy that maximizes relevance and diversity, which achieve higher scores. This gap reflects MLVU’s focus on fine-grained temporal grounding and multi-event reasoning, where heuristics can sometimes capture domain-specific cues more effectively. Overall, the results demonstrate that even without the K Head, the Rank Head alone can reliably prioritize important frames and deliver consistent gains across multiple VLMs, achieving state-of-the-art performance on most benchmarks.
RQ 3: How much can FrameOracle reduce computational cost while preserving accuracy?
We take LLaVA-Video-7B with 16 input frames as the baseline and report per-GPU, per-sample averages. FrameOracle reduces the input from 16 to 10.4 frames, cutting the VLM cost from 184.38 to 109.11 TFLOPs and the end-to-end total from 184.38 to 110.98 TFLOPs ($39.8\%$). It also lowers latency from 0.615 to 0.363 seconds ($41.0\%$) and reduces tokens from 11,644.0 to 7,581.6, while maintaining accuracy. With a larger candidate pool, FrameOracle reduces 64 frames to 13.9, improving accuracy by +1.4+1.4 while still lowering total compute to 167.67 TFLOPs ($9.1\%$), tokens to 10,133.1 ($13.0\%$), and latency to 0.556 seconds ($9.6\%$). These results reveal a clear trade-off: smaller pools yield larger efficiency gains with no loss in accuracy, while larger pools deliver accuracy gains with moderate compute savings. Although reported on LLaVA-Video-7B, computation is dominated by the backbone VLM, so similar efficiency–accuracy trade-offs are expected for other $\sim$7B-scale models, with only minor variation due to architectural details.
RQ 4: Are all training-stage components essential for FrameOracle’s performance?
We conduct ablations over the four training stages to evaluate the contribution of each stage, using Qwen2.5-VL-3B as the backbone VLM (Table 4). The baseline (first row) randomly selects 16 frames from 32 uniformly sampled candidates. Stage 1 (text–visual alignment) underperforms the baseline, though it provides a foundational starting point. Stages 2 (Rank Head) and 3 (K Head) yield clear performance improvements, and the full model with Stage 4 (fine-tuning on FrameOracle-41K) delivers further gains over Stages 2 and 3 on most benchmarks, with only a slight decline on Video-MME. Moreover, the average number of retained frames decreases from 21.8 to 20.9, showing that FrameOracle-41K supervision stabilizes performance while enabling higher accuracy with fewer frames. These results demonstrate that ground-truth supervision from FrameOracle-41K is essential for refining both frame importance scoring and the prediction of the number of frames, establishing it as a valuable resource for adaptive frame selection.
@@
Table 4: Four-stage training of FrameOracle, evaluated on Qwen2.5-VL-3B. Stages are added progressively to assess their impact. The baseline (first row) randomly selects 16 of 32 frames. Bold numbers indicate best performance. LVB = LongVideoBench validation set.
Model Frames NExTQA Perception LVB Video-MME EgoSchema MLVU
OE_val OE_test MC
[…]
@@ 6 Conclusion
We propose FrameOracle, a lightweight, plug-and-play frame selector that adaptively determines which frames to retain and how many are needed. To facilitate training, we introduce FrameOracle-41K, a large-scale VideoQA dataset with 40,992 examples, and the first to provide keyframe annotations specifying the minimal frames required to answer each question. Extensive experiments show that FrameOracle improves the performance of diverse VLM backbones without co-training, reducing FLOPs, latency, and the number of tokens, while also outperforming state-of-the-art keyframe selection methods. Future work will explore relaxing the fixed-length input setting to support variable-sized frame sets.
Reproducibility Statement
To support reproducibility, we provide details on both the model and dataset. FrameOracle’s design, including its learning objective, selection policy, and staged curriculum, is described in Section 4, with training procedures and hyperparameters in Appendix A (Figure 4; Table 5). FrameOracle-41K construction, including agent-based mining, verification, and dataset format, is covered in Section 3 and Appendix B. Evaluation settings, including benchmarks, backbones, and metrics, are in Section 5.1. These sections provide all the information needed to reproduce our results.
Ethics Statement
FrameOracle-41K is created from source videos collected from the internet, which may contain images of individuals and reflect societal biases present in online content. Our data processing pipeline does not involve identifying or profiling any individuals. The data is used solely for developing our video understanding model. We release the dataset strictly for non-commercial, academic research purposes and caution future users to be aware of potential inherent biases in the data.

### v2

@@ 4.2 Training
Jointly learning which frames to select and how many to retain is challenging, as optimizing frame ranking and the predicted subset size simultaneously can destabilize training. To address this, we adopt a staged training strategy that gradually refines the selection policy $\Pi_{\theta}$: first learning robust frame and text representations, then frame ranking, and finally predicting the number of frames to keep. This approach enables effective reasoning over pre-sampled frame sets (e.g., 16 or 64 frames). Training leverages four widely used public VideoQA datasets covering clips from roughly 10 seconds to 15 minutes (details in Appendix A). The necessity of this multi-stage curriculum is further validated through targeted ablation studies, detailed in Appendix F.1.
Stage 1: Text-Visual Alignment. This stage trains FrameOracle to produce robust cross-modal representations by aligning the textual query with the candidate frame set $V_{C}$. We use the pre-trained SigLIP model (Zhai et al., 2023) as a teacher, treating its similarity scores for each query-frame pair as target relevance signals. The frame and text feature projectors, along with the cross-modal Transformer encoder, are trained using a RankNet loss (Burges et al., 2005) to match the relative ordering of SigLIP similarities. Concretely, for frames $i$ and $j$ with SigLIP scores $s_{i}$ and $s_{j}$, and predicted scores $y_{i}$ and $y_{j}$, we define the pairwise label $t_{ij}=\mathrm{sign}(s_{i}-s_{j})$. The loss is
@@
where $t_{ij}=0$ (ties) do not contribute to gradients. The K Head is frozen during this stage, ensuring that learning focuses solely on frame ranking.
Stage 2: Rank Head Optimization. In this stage, the Rank Head is trained to identify the most salient frames in the candidate set $V_{C}$. Unlike Stage 1, which uses frame-level SigLIP scores independently, we leverage the downstream VLM’s loss to capture frame importance in context, including temporal dependencies. For each frame $f_{i}\in V_{C}$, we adopt a leave-one-out (LOO) approach: $f_{i}$ is removed from the input set, and the resulting change in the VLM’s loss indicates its importance. These LOO-based scores serve as soft targets, and the Rank Head is trained with a RankNet loss to predict them. The Transformer encoder and feature projectors are fine-tuned at a smaller learning rate for stability, while the K Head remains frozen.
Stage 3: K Head Optimization. In this stage, the K Head learns to predict the number of frames to select, $K$. The Rank Head is frozen, while the Transformer encoder and feature projectors are fine-tuned with a very small learning rate for slight adaptation. For each sample, we evaluate the downstream VLM on subsets of top-$k$ frames from $V_{C}$, ranked by the frozen Rank Head, for $k\in\{1,\dots,N\}$. The target number of frames is chosen as
$k^{*}\;=\;\arg\min_{k\in\{1,\dots,N\}}\Big(\mathrm{zscore}(\mathcal{L}_{\text{task}}(k))+\lambda_{k}\,k\Big),$ (2)
where the linear penalty $\lambda_{k}k$ balances accuracy and frame cost. The K Head predicts a categorical distribution $p_{\theta}(k)$ over possible $k$ values and is trained with a combination of the Expected Value Objective and a soft-target classification loss:
@@
regresses the expected frame count to $k^{*}$, and $\mathcal{L}_{\text{class}}$ aligns $p_{\theta}$ with a Gaussian-shaped soft target centered at $k^{*}$.
Stage 4: Supervised Fine-tuning with Keyframe Annotations. In the final stage, FrameOracle is fine-tuned on FrameOracle-41K, which provides verified annotations for both keyframe indices and the number of frames, $K$. The Rank Head is trained to align its predictions with the annotated keyframes, while the K Head is jointly trained to predict the annotated $K$ values. Unlike the previous stages, which rely on weak or proxy signals, this direct supervision consolidates the selection policy, ensuring that FrameOracle accurately identifies both what to see and how much to see for each query.
@@ 5.1 Experiment Settings
Benchmarks. We evaluate FrameOracle on six widely adopted video benchmarks, which can be divided into long-video and short-video understanding tasks. For long-video understanding, we include EgoSchema (Mangalam et al., 2023) (all videos 3 min), LongVideoBench (Wu et al., 2024) (average $\sim$12 min), MLVU (Zhou et al., 2025) (average $\sim$12 min), and Video-MME (Fu et al., 2025) (average $\sim$17 min), all of which require reasoning over extended temporal contexts ranging from minutes to hours. These datasets emphasize challenges such as cross-event reasoning, global consistency, and temporal grounding across lengthy sequences. For short-video understanding, we evaluate on NExTQA (Xiao et al., 2021) and Perception (pătrăucean2023perceptiontest), which contain clips typically under one minute and focus on fine-grained event recognition, local temporal relations, and reasoning within concise videos. Evaluation follows the LMMs-Eval library (Zhang et al., 2025b), and we report accuracy across all benchmarks.
@@
VideoChat2-7B (Li et al., 2024) 16 - - - - 39.3 39.5 63.6 44.5 -
@@
[…]
Table 1: FrameOracle vs. SOTA VLMs. “Frames” shows $M\rightarrow\bar{K}$: FrameOracle starts from $M$ uniformly sampled frames and reduces to an average of $\bar{K}$ frames. Highlighted rows correspond to larger frame inputs. LVB = LongVideoBench validation set. For each backbone model, statistical significance is evaluated for the best-performing FrameOracle variant. 95% confidence intervals (CIs) are computed via paired bootstrap resampling over questions (10,000 iterations, percentile method). Improvements are considered statistically significant at $\alpha=0.05$ if the CI excludes zero. § denotes non-significant changes (95% CI includes zero).
@@
[5pt/3pt] Qwen2.5-VL-7B (Bai et al., 2025) 8 69.2 53.6 54.1 56.8 54.5
[…]
Table 2: FrameOracle vs. SOTA keyframe selection methods. NExTQA reports MCQ. Methods using more frames or larger LLMs are shown in gray. LVB = LongVideoBench validation set. For fair comparison under a fixed budget, we disable FrameOracle’s K Head and use only the Rank Head to select the top-8 frames from uniformly sampled candidates. For all backbone models, FrameOracle achieves statistically significant improvements over the corresponding baselines (paired bootstrap, $\alpha=0.05$).
5.2 Results and Analysis
Comparison with State-of-the-Art and FrameOracle-Enhanced Models. Table 5.1 compares FrameOracle under two evaluation settings: (1) standalone performance of four state-of-the-art (SOTA) VLM baselines, and (2) FrameOracle integrated as a frame selector with five diverse VLM backbones. For each backbone integrated with FrameOracle, we evaluate two configurations, using candidate pools of 16 and 64 uniformly sampled frames. The Qwen-VL series internally merges every two adjacent frames into a single visual representation. To ensure a fair comparison with models that process raw frames, we report Qwen baselines with 32 and 128 input frames.
With a 16-frame candidate pool, FrameOracle preserves accuracy across all benchmarks while reducing the number of frames processed by the downstream VLM by approximately 35%. When scaling the candidate pool to 64 frames, FrameOracle benefits from denser temporal coverage and adaptively selects more informative subsets, consistently improving accuracy over the corresponding baselines by an average of 1.5% while still reducing the effective frame count by about 15%. These results indicate that larger candidate pools allow FrameOracle to better exploit temporal redundancy, yielding stronger accuracy–efficiency trade-offs.
This scaling behavior is particularly evident in long-video settings. For example, starting from 128 candidate frames, Qwen3-VL-8B equipped with FrameOracle outperforms Video-XL-7B, which processes 256 frames, and even Video-XL-2-8B, which operates on approximately 10k frames, on LongVideoBench (average video length $\sim$12 min) and Video-MME (average video length $\sim$17 min). Since FrameOracle is trained independently and applied in a fully plug-and-play manner, without co-training or backbone-specific adaptation, these results demonstrate its strong generalization across model architectures.
One might expect that providing more frames improves performance by offering additional visual evidence. However, Table 5.1 shows that using more frames often fails to help and can even reduce accuracy. This observation aligns with recent findings that long-video reasoning is inherently sparse, with only a small subset of frames being truly relevant (Park et al., 2024). Additional frames primarily introduce redundancy and noise, which can dilute cross-modal attention and lead to diminishing returns (Li et al., 2023).
In contrast, when FrameOracle selects a smaller but more informative subset of frames, performance improves, particularly on open-ended benchmarks. For example, on LLaVA-OneVision-7B, FrameOracle reduces the input from 16 frames to about 10 frames on average, while improving NExTQA performance (OE_val: $14.6\rightarrow 16.1$, OE_test: $16.7\rightarrow 17.8$) and EgoSchema ($60.8\rightarrow 62.4$). Qualitative examples in Appendix I further illustrate that FrameOracle isolates the critical visual evidence, yielding correct answers where naive higher-frame sampling fails.
Crucially, these improvements do not arise from simply reducing the number of input frames. As shown in our ablation study (Appendix F.3), uniformly sampling 10 frames from the same 16-frame input substantially degrades performance (49.8% $\rightarrow$ 46.3% on LLaVA-OneVision-7B). In contrast, FrameOracle improves accuracy to 50.3% with the same frame budget. This demonstrates that the gains stem from selecting semantically relevant frames, rather than merely reducing the number of visual tokens.
@@ 5.2 Results and Analysis
We compare FrameOracle with existing keyframe selection methods, including (1) jointly trained approaches and (2) plug-and-play selectors applied to open-source VLMs (Table 5.1). Because jointly trained methods are not directly transferable, we focus on plug-and-play comparisons under a unified 8-frame evaluation setting. For fairness, we disable FrameOracle’s K Head and use only the Rank Head to select the top-8 frames from uniformly sampled candidates.
FrameOracle performs strongly across multiple backbones and benchmarks. On LLaVA-OneVision-7B and LLaVA-Video-7B, it consistently outperforms existing plug-and-play selectors such as Frame-Voyager (Yu et al., 2025), BOLT (Liu et al., 2025), and KFC (Fang et al., 2025) on NExTQA, LongVideoBench, Video-MME, and EgoSchema. On the Qwen2.5-VL-7B backbone, FrameOracle also compares favorably against recent strong baselines, outperforming ViaRL (Xu et al., 2025) and matching or exceeding K-frames (Yao et al., 2025), while operating with a smaller candidate pool (128 vs. 256 frames).
On MLVU, FrameOracle improves over the base VLMs but does not consistently outperform all SOTA methods when using a 64-frame candidate pool. MLVU includes very long videos, some exceeding 2 hours, emphasizing broad temporal coverage. Several SOTA methods address this challenge by operating with substantially higher temporal coverage, using 128 or 256 frames, or sampling at fixed rates such as 1 FPS across the entire video. To explore how FrameOracle behaves under increased coverage, we evaluate a variant that divides the video into multiple temporal segments and applies FrameOracle independently within each segment, effectively increasing coverage while keeping the final frame budget fixed. Results in Appendix F.4 show that this strategy leads to improved performance on MLVU.
Overall, these results show that the Rank Head alone can reliably prioritize informative frames and achieve state-of-the-art performance across most benchmarks, even without adaptive frame-count prediction.
@@
RQ 3: How effectively can FrameOracle reduce computation while maintaining accuracy? We take LLaVA-Video-7B with 16 input frames as the baseline and report per-GPU, per-sample averages. FrameOracle reduces the input from 16 to 10.4 frames, cutting the VLM cost from 184.38 to 109.11 TFLOPs and the end-to-end total to 110.98 TFLOPs ($39.8\%$). It also lowers latency from 0.615 to 0.363 seconds ($41.0\%$) and reduces tokens from 11,644.0 to 7,581.6, while maintaining accuracy. With a larger candidate pool, FrameOracle reduces 64 frames to 13.9, improving accuracy by +1.4+1.4 while still lowering total compute to 167.67 TFLOPs ($9.1\%$), tokens to 10,133.1 ($13.0\%$), and latency to 0.556 seconds ($9.6\%$). These results highlight a clear trade-off: smaller candidate pools yield larger efficiency gains without harming accuracy, while larger pools enable more significant accuracy improvements with moderate compute savings. Notably, roughly 90% of the total computation comes from the backbone VLM, with FrameOracle contributing only about 10% of the cost. As shown in Table 3, moving from the 16 → 10.4 to the 64 → 13.9 setting increases total FLOPs almost entirely due to the VLM. While the exact trade-offs vary, the observed efficiency–accuracy patterns are likely to extend across models of different sizes and architectures, as they primarily stem from the relative cost distribution between the lightweight selector and the backbone VLM.
@@ 6 Conclusion
We propose FrameOracle, a lightweight, plug-and-play frame selector that adaptively determines both which frames to select and how many are needed from a given candidate pool. To support training, we introduce FrameOracle-41K, a large-scale VideoQA dataset with 40,992 examples and the first to provide question-conditioned keyframe annotations that specify the minimal frames required to answer each query. Extensive experiments demonstrate that FrameOracle consistently improves a wide range of VLM backbones without co-training, while reducing FLOPs, inference latency, and visual token usage, and outperforming state-of-the-art keyframe selection methods. While FrameOracle-41K provides high-quality, large-scale supervision for adaptive frame selection, expanding the dataset with more diverse videos and question types could further strengthen frame-level supervision and enable training FrameOracle primarily from strong keyframe annotations, potentially simplifying the current multi-stage training pipeline.
Impact Statement
This paper presents work whose goal is to advance the field of machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.